OpenStudy (anonymous):

which of the following is the equation of a line that passes through the points (1, 6) and (2, 1) a. y=-5x+1 b.y=2x+1 c.y=5x-1 d.y=-5x+11

OpenStudy (nincompoop):

do you know how to solve for slope?

OpenStudy (anonymous):

rise/run right?

OpenStudy (nincompoop):

yeahhhhh SOLVE FOR THE SLOPE FIRST

OpenStudy (anonymous):

use this (y-y1)/(x-x1)=(y2-y1)/(x2-x1) you have no need to find slope 1st for this eq..just plug in values and solve..

OpenStudy (nincompoop):

@zaibali.qasmi I got this, you can try helping other kids

OpenStudy (nincompoop):

and don't believe him

OpenStudy (nincompoop):

let us do this step by step solve for the slope first

OpenStudy (anonymous):

oh so that's not right...I started plugging it in

OpenStudy (nincompoop):

it is not correct

OpenStudy (anonymous):

ho can say its not corect..???if i prove it then..????

OpenStudy (nincompoop):

he is using similar techniques but all in one equation

OpenStudy (nincompoop):

you don't need to learn that for now because it will only confuse you

OpenStudy (anonymous):

okay thank you, So what would I plug in and where would I plug it in to find slope? im sorry I'm teaching this to myself I'm a bit confused ^^"

OpenStudy (nincompoop):

look at your given points (1, 6) and (2, 1) assign x1, y1 and x2, y2

OpenStudy (anonymous):

okay I know that

OpenStudy (nincompoop):

@zaibali.qasmi chill just help other kids you're doing this wrong math is not about plugging in

OpenStudy (anonymous):

@nincompoop @ghosty if i prove it by using that one which suggest then..??? its option D by hat eq as i suggest..and am going to prove here,,,,see next..

OpenStudy (nincompoop):

someone's pride is hurt I can see

OpenStudy (anonymous):

d

OpenStudy (nincompoop):

@sochittasal try giving the answer next time and I will have you banned for cheating

OpenStudy (anonymous):

\[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\] \[\frac{ y-6 }{ x-1 }=\frac{ 1-6 }{ 2-1 }\] \[\frac{ y-6 }{ x-1 }=\frac{ -5 }{ 1 }\] now cross multiplication gives... (y-6)*1=-5(x-1) => y-6=-5x+5 =>y=-5x+5+6 =>y=-5x+11 see its exactly option D...

OpenStudy (anonymous):

thank you guys for helping out

OpenStudy (nincompoop):

I think that method is longer and it does not help with the concept

OpenStudy (anonymous):

i appreciate it very much

OpenStudy (nincompoop):

I bet you, she will ask another problem of the same concept because it wasn't taught properly

OpenStudy (anonymous):

@nincompoop ...its not pride ..its selfrespect...if you dont want someone else interaction then just going on everything..why you call that your method is wrong..???

OpenStudy (anonymous):

@see above what she say..i appriciate it much..now lolllzzz.

OpenStudy (nincompoop):

because it is wrong and unintuitive the reason they are asking for help is because they do not understand the concept and you're just asking people to plug numbers in math is not about plugging things in

OpenStudy (nincompoop):

you have to go in their level of understanding and not assume that they know what that equation you gave

OpenStudy (anonymous):

when there is a way to solve question ...in easy way..then y you are taking asker to a hard one..??

OpenStudy (nincompoop):

it is not more difficult, in fact yours require more step that is cross multiplication

OpenStudy (nincompoop):

people here get simple multiplication wrong